# FLIP: Towards Fine-grained Alignment between ID-based Models and Pretrained Language Models for CTR Prediction

### v3

@@ IV-A Overview of FLIP
Hereinafter, we omit the detailed structure of PLMs and ID-based models, since FLIP serves as a model-agnostic framework and is compatible for various backbone models.
IV-B Modality Transformation
@@
Standard PLMs take sequences of words as inputs [19, 47]. Consequently, the modality transformation aims to convert the tabular data $\textbf{x}_{i}^{tab}$ into textual data $\textbf{x}_{i}^{text}$ via hard prompt templates. Previous works [25, 21, 70] have suggested that sophisticated templates (e.g., with more vivid descriptions) for textual data construction might mislead the language model and make it fails to grasp the key information in the texts. Hence, we adopt the following simple yet effective transformation template:
$t_{i,f}$ $=[m_{f},\text{``}is\text{''},v_{i,f},\text{``}.\text{''}],\;f\in\{1,\cdots,F\}$ (3)
$\textbf{x}^{text}_{i}$ $=[t_{i,1},t_{i,2},\cdots,t_{i,F}],$
@@
where $m_{f}$ is the name of $f$-th field (e.g., gender), $v_{i,f}$ denotes the feature value of $f$-th field for input $x_{i}^{tab}$ (e.g., female), and []delimited-[][$\cdot$][ ⋅ ] indicates the concatenation operator.
@@
An illustrative example is given in Figure 3, where we first construct a descriptive sentence $t_{i,f}$ for each feature field, and then concatenate them to obtain the final textual data $\textbf{x}_{i}^{text}$. In this way, we preserve the semantic knowledge of both field names and feature values with minimal preprocessing and no information loss.
IV-C Modality Alignment Pretraining
After obtaining the paired text-tabular data $(\textbf{x}_{i}^{text},\textbf{x}_{i}^{tab})$ for the same original raw input, we design a novel modality alignment pretraining algorithm to achieve both instance-level and feature-level alignments for PLMs and ID-based models.
@@ IV-C Modality Alignment Pretraining
As shown in Figure 4, we first perform field-level data masking to obtain the corrupted version of input pair, i.e., $(\hat{\textbf{x}}_{i}^{text},\hat{\textbf{x}}_{i}^{tab})$. Then, the PLM $h_{\text{PLM}}$ and ID-based model $h_{\text{ID}}$ encode the input data to obtain the dense representations $(\textbf{w}_{i},\hat{\textbf{w}}_{i})$ and $(\textbf{v}_{i},\hat{\textbf{v}}_{i})$ for textual and tabular modalities, respectively. Next, we apply three different pretraining objectives over the dense representations for loss computation and model optimization:
@@
  • Joint reconstruction for Masked Language Modeling (MLM) requires the model to recover the original tokens from the corrupted textual context with the help of complete tabular data.
@@
  • Joint reconstruction for Masked Tabular Modeling (MTM) requires the model to recover the original feature IDs from the corrupted tabular context with the help of complete textual data.
@@
While ICL only provides coarse-grained self-supervisions, the joint reconstruction for MLM/MTM casts a dual pretraining task with fine-grained self-supervised signals. That is, each single model (i.e., PLM or ID-based model) has to seek and exploit the useful knowledge underlying the representation from the other modality to accomplish the recovery task, resulting in fine-grained feature-level cross-modal alignments.
IV-C1 Field-level Data Masking
We introduce the field-level data masking strategy for textual and tabular data, respectively.
For textual data, we propose to mask tokens at the field level. We first randomly select a certain ratio $r_{text}$ of the feature fields to be corrupted, and then mask all the consecutive tokens that constitute the entire text of the corresponding feature value. Note that this is quite different from the original random token-level masking strategy for common language models (e.g., BERT [19]), which is not a good choice for FLIP. The hard template we adopt has fixed patterns and structures, i.e., $\textbf{x}_{i}^{text}$ is simply a series of “what is what” sentences, and the descriptive text for each feature is also fixed. Hence, a random token-level masking strategy might lead to a too-easy pretraining task, where the model could suffer from the early degeneration problem and lose the generalization ability [71]. Besides, we denote the index set of masked tokens as $\mathcal{I}^{text}$.
As an example for textual data masking, suppose the tokenization of sentence for occupation field is [“occupation”, “is”, “college”, “student”]. The outcome of field-level masking should be [“occupation”, “is”, [MASK], [MASK]]. If the token-level masking is adopted, the outcome might be [[MASK], “is”, “college”, “student”] or [“occupation”, “is”, “college”, [MASK]]. Obviously, these masked tokens can be easily inferred solely based on the textual context, thus limiting the cross-modal interaction for joint reconstruction.
For tabular data, following previous works [7, 72], we uniformly sample a certain ratio $r_{tab}$ of the fields, and then replace the corresponding feature with an additional <MASK> feature, which is set as a special feature in the embedding table of ID-based model. The <MASK> feature is not field-specific, but shared by all feature fields to prevent introducing prior knowledge about the masked field [7]. Similarly, we denote the index set of masked fields as $\mathcal{I}^{tab}$.
@@ IV-C2 Data Encoding
$\textbf{w}_{i}$ $=h_{\text{PLM}}(\textbf{x}_{i}^{text}),\;\;\;\hat{\textbf{w}}_{i}=h_{\text{PLM}}(\hat{\textbf{x}}_{i}^{text}),$ (4)
@@
Here, $\textbf{w}_{i}=[w_{i,l}]_{l=1}^{L}\in\mathbb{R}^{L\times D_{text}}$ is the set of hidden states from the last layer of language model, where $L$ is the number of tokens for $\textbf{x}_{i}^{text}$ and $D_{text}$ is the hidden size of language model. Note that $w_{i,1}$ is the [CLS] token vector that represents the overall textual input. $\hat{\textbf{w}}_{i}=[\hat{w}_{i,l}]_{l=1}^{L}\in\mathbb{R}^{L\times D_{text}}$ also satisfies the notations above. $\textbf{v}_{i},\hat{\textbf{v}}_{i}\in\mathbb{R}^{D_{tab}}$ are the representations produced by ID-based model.
IV-C3 Joint Reconstruction for Masked Language Modeling
As shown in Figure 4, the MLM module takes the text-tabular pair $(\hat{\textbf{w}}_{i},\textbf{v}_{i})$ as input, and attempts to recover the masked tokens. We denote the index set of masked tokens as $\mathcal{I}^{text}$. For each masked token with index $l\in\mathcal{I}^{text}$, we concatenate the corresponding token vector $\hat{w}_{i,l}$ with the “reference answer” $\textbf{v}_{i}$, and then feed them through the prediction layer to obtain the estimated distribution:
$q_{i,l}=g_{\text{PLM}}([\hat{w}_{i,l}\oplus\textbf{v}_{i}])\in\mathbb{R}^{V},$ (5)
where $\oplus$ is the concatenation operation, $V$ is the vocabulary size, and $g_{\text{PLM}}$ is a two-layer MLP similar to the decoder of language model. Finally, similar to common masked language modeling [19], we leverage the cross-entropy loss for pretraining optimization:
$\mathcal{L}^{MLM}_{i}=\frac{1}{|\mathcal{I}^{text}|}\sum\nolimits_{l\in\mathcal{I}^{text}}\operatorname{CrossEntropy}\left(q_{i,l},\ x^{text}_{i,l}\right),$ (6)
@@ IV-C3 Joint Reconstruction for Masked Language Modeling
IV-C4 Joint Reconstruction for Masked Tabular Modeling
@@
$\textbf{u}_{i}=\text{Softmax}\left(\frac{\widehat{\textbf{v}}_{i}\textbf{Q}\textbf{w}_{i}^{T}}{\sqrt{D_{text}}}\right)\textbf{w}_{i},\;\;\textbf{u}_{i}\in\mathbb{R}^{L}$ (7)
where $\textbf{Q}\in\mathbb{R}^{D_{tab}\times D_{text}}$ is the trainable cross-modal attention matrix, and $\sqrt{D_{text}}$ is the scaling factor [73].
@@
$\textbf{c}_{i,f}$ $=g_{\text{ID}}^{(f)}(\textbf{u}_{i}),\quad\textbf{c}_{i,f}\in\mathbb{R}^{M},$ (8)
@@
$\mathcal{L}_{i}^{MTM}=\frac{1}{|\mathcal{I}^{tab}|}\sum\nolimits_{f\in\mathcal{I}^{tab}}\operatorname{CrossEntropy}(p_{i,f},\ x^{tab}_{i,f}),$ (9)
@@
However, the loss above is actually impractical and inefficient since it has to calculate the softmax function over the entire feature space in Eq. 8, where $M$ is usually at million level for real-world applications. To this end, we adopt noise contrastive estimation (NCE) [74, 75, 7]. NCE transforms a multi-class classification task into a binary classification task, where the model is required to distinguish positive features (i.e., masked feature) from noise features. Specifically, for each masked feature $x_{i,f}^{tab}$ of $f$-th field, we sample $K$ noise features from the entire feature space according to their frequency distribution in the training set. Then, we utilize the binary cross-entropy (BCE) loss for MTM optmization:
$\mathcal{L}_{i}^{MTM}=-\frac{1}{|\mathcal{I}^{tab}|}\sum_{f\in\mathcal{I}^{tab}}(\log\sigma(c_{i,f,t})+\sum_{k=1}^{K}\log(1-\sigma(c_{i,f,k}))))$ (10)
@@
IV-C5 Instance-level Contrastive Learning (ICL)
In addition to the feature-level alignment brought by the joint reconstruction, we further introduce contrastive learning to explicitly establish instance-level alignment between two modalities. The contrastive objective draws the representations of matched text-tabular pairs together and pushes apart those non-matched pairs [34, 76].
In the ICL module, we leverage the [CLS] token vector $w_{i,1}$ to represent the textual input $\textbf{x}_{i}^{text}$. For dimensional consistency, we first employ two separate linear layers followed by a normalization layer to project the [CLS] token representation $w_{i,1}$ and tabular representation $\textbf{v}_{i}$ into $d$-dimensional vectors, $z^{text}_{i}$ and $z^{tab}_{i}$, respectively. $d$ is set to 128 in this paper. Next, we adopt InfoNCE [77] to compute the ICL loss:
$\mathcal{L}_{text2tab}^{ICL}$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\left(\frac{\text{exp}(\text{sim}(z^{text}_{i},z^{tab}_{i})/\tau)}{\sum_{j}\text{exp}(\text{sim}(z^{text}_{i},z^{tab}_{j})/\tau)}\right),$ (11)
$\mathcal{L}_{tab2text}^{ICL}$ $=-\frac{1}{B}\sum_{i=1}^{B}\log\left(\frac{\text{exp}(\text{sim}(z^{tab}_{i},z^{text}_{i})/\tau)}{\sum_{j}\text{exp}(\text{sim}(z^{tab}_{i},z^{text}_{j})/\tau)}\right),$ (12)
$\mathcal{L}^{ICL}$ $=\frac{1}{2}(\mathcal{L}_{text2tab}^{ICL}+\mathcal{L}_{tab2text}^{ICL}),$ (13)
@@ IV-C5 Instance-level Contrastive Learning (ICL)
$\mathcal{L}^{pretrain}=\frac{1}{B}\sum\nolimits_{i=1}^{B}\left(\mathcal{L}^{MLM}_{i}+\mathcal{L}^{MTM}_{i}\right)+\mathcal{L}^{ICL}.$ (14)
[…]
IV-D Supervised Finetuning
After the pretraining stage, the PLM and ID-based model have received better parameter initialization with fine-grained cross-modal alignments. As depicted in Figure 5(a), in this stage, FLIP finetunes the two models jointly on the downstream CTR prediction task with supervised click signal to achieve superior performance.
FLIP places a randomly initialized linear layer on the ID-based model, and another layer upon the PLM, so that these two model can output the estimated probability $\hat{y}_{i}^{\text{ID}}$ and $\hat{y}_{i}^{\text{PLM}}$ respectively.
$\hat{y}^{\text{ID}}_{i}$ $=\sigma(\operatorname{Linear}_{\text{ID}}(\textbf{v}_{i})),$ (15)
@@ IV-D Supervised Finetuning
$\hat{y}_{i}=\sigma\left(\alpha\times\operatorname{Linear}_{\text{ID}}(\textbf{v}_{i})+(1-\alpha)\times\operatorname{Linear}_{\text{PLM}}(\textbf{w}_{i})\right),$ (16)
where $\alpha\in$ is a learnable parameter to balance the outcomes from two different modalities. To avoid performance collapse mentioned in previous works [78, 79], we apply BCE objectives over the jointly estimated click probability $\hat{y}$, as well as the solely estimated click probability $\hat{y}_{i}^{\text{ID}}$ and $\hat{y}_{i}^{\text{PLM}}$ for model optimization:
$\mathcal{L}^{finetune}=\mathcal{L}_{BCE}(y,\hat{y})+\mathcal{L}_{BCE}(y,\hat{y}^{\text{ID}})+\mathcal{L}_{BCE}(y,\hat{y}^{\text{PLM}})$ (17)
@@
$\mathcal{L}_{PLM}$ $=\mathcal{L}_{BCE}(y,\hat{y}^{\text{PLM}}).$ (19)
[…]
V Experiment
In this section, we aim to evaluate the performance and effectiveness of the proposed method. Specifically, we conduct extensive experiments to answer the following research questions:
  • RQ1
    Does our proposed FLIP outperform existing state-of-the-art baseline models for CTR prediction?
  • RQ2
    Does FLIP have high compatibility with various ID-based models and PLMs?
  • RQ3
    How do the main components (e.g., MLM or MTM) of FLIP affect the performance?
  • RQ4
    How do different hyperparameter settings (e.g., the mask ratio) of FLIP impact the performance?
  • RQ5
    Does FLIP really establish the feature-level alignment?
  • RQ6
    Does FLIP promote the feature embedding learning, and alleviate the representation degeneration problem?
V-A Experiment Setup
V-A1 Datasets
We conduct extensive experiments on three real-world public datasets. All of the selected datasets contain user and item information. The information is unencrypted original text, thus preserving the real semantic information.
  • MovieLens-1M [80] is a movie recommendation dataset with user-movie ratings ranging from 1 to 5. Following the previous work [17, 25], We consider samples with ratings greater than 3 as positive, samples with ratings less than 3 as negative, and remove samples with ratings equal to 3 (i.e., neutral).
  • BookCrossing [81] is a book recommendation dataset and possesses user-book ratings ranging from 0 to 10. We consider samples with scores greater than 5 as positive, and the rest as negative.
  • GoodReads [82, 83] is a book recommendation dataset which contains user-book ratings ranging from 1 to 5. We take samples with ratings greater than 3 as positive, and the rest as negative.
The statistics of the processed datasets are shown in Table I. Following previous works [25, 84], we sort all samples in chronological order and take the first 90% samples as the training set and the remaining as the testing set. The training set for the pretraining stage and finetuning stage are the same [7]. All our experimental results are obtained on the testing set.
TABLE I: Statistics of processed datasets.

### v4

@@ 3.1. Overview of FLIP
Hereinafter, we omit the detailed structure of PLMs and ID-based models, since FLIP serves as a model-agnostic framework and is compatible with various backbone models.
3.2. Modality Transformation
@@
Standard PLMs take sequences of words as inputs (Devlin et al., 2018; Radford et al., 2018). Consequently, the modality transformation aims to convert the tabular data $\textbf{x}_{i}^{tab}$ into textual data $\textbf{x}_{i}^{text}$ via hard prompt templates. Previous works (Li et al., 2023a; Liu et al., 2022; Hegselmann et al., 2023) have suggested that sophisticated templates (e.g., with more vivid descriptions) for textual data construction might mislead the PLM and make it fail to grasp the key information in the texts. Hence, we adopt the following simple yet effective transformation template:
(3) $t_{i,f}$ $=[m_{f}\oplus\text{``}is\text{''}\oplus v_{i,f}\oplus\text{``}.\text{''}],\;f\in\{1,\cdots,F\}$
$\textbf{x}^{text}_{i}$ $=[t_{i,1}\oplus t_{i,2}\oplus\cdots\oplus t_{i,F}],$
@@
where $m_{f}$ is the name of $f$-th field (e.g., gender), $v_{i,f}$ denotes the feature value of $f$-th field for input $x_{i}^{tab}$ (e.g., female), and $\oplus$ indicates the concatenation operator.
@@
An illustrative example is given in Figure 2 (Stage 1), where we first construct a descriptive sentence $t_{i,f}$ for each feature field, and then concatenate them to obtain the final textual data. In this way, we preserve the semantic knowledge of both field names and feature values with minimal preprocessing and no information loss. Both tabular data and textual data can be considered to contain almost the same information of raw data, albeit in different modalities.
3.3. Modality Alignment Pretraining
@@
As shown in Figure 2 (Stage 2), after obtaining the paired text-tabular data $(\textbf{x}_{i}^{text},\textbf{x}_{i}^{tab})$ from the same raw input, we first perform field-level data masking to obtain the corrupted version of input pair, $(\hat{\textbf{x}}_{i}^{text},\hat{\textbf{x}}_{i}^{tab})$. Then, the PLM $h_{\text{PLM}}$ and ID-based model $h_{\text{ID}}$ encode the input pair to obtain the dense representations $(\textbf{w}_{i},\hat{\textbf{w}}_{i})$ and $(\textbf{v}_{i},\hat{\textbf{v}}_{i})$ for textual and tabular modalities, respectively. Next, we apply three different pretraining objectives to achieve both feature-level and instance-level alignments between PLMs and ID-based models:
@@
  • Feature-level Masked Language Modeling (MLM) requires the model to recover the original tokens from the corrupted textual context with the help of complete tabular data.
@@
  • Feature-level Masked Tabular Modeling (MTM) requires the model to recover the original feature IDs from the corrupted tabular context with the help of complete textual data.
@@
The jointly masked modality modeling task learns fine-grained modality interactions by the way of mask-and-predict, while ICL aligns the modalities from the perspective of the global consistency.
3.3.1. Field-level Data Masking
We first perform the field-level data masking strategy for textual and tabular data, respectively.
For textual data, we propose to mask tokens at the field level. We first randomly select a certain ratio $r_{text}$ of the feature fields to be corrupted, and then mask all the consecutive tokens that constitute the entire text of the corresponding feature value. We denote the index set of masked tokens as $\mathcal{I}^{text}$. Note that this is quite different from the random token-level masking strategy for common language models (e.g., BERT (Devlin et al., 2018)).
For example, suppose the sentence tokenization from occupation field is [“occupation”, “is”, “college”, “student”]. The outcome of field-level masking should be [“occupation”, “is”, [MASK], [MASK]]. But the outcome of token-level masking might be [[MASK], “is”, “college”, “student”] or [“occupation”, “is”, “college”, [MASK]]. Obviously, token-level masked tokens can be easily inferred solely based on the textual context, thus limiting the cross-modal interactions and losing the generalization ability (Liu et al., 2021).
For tabular data, following previous works (Lin et al., 2023b; Wang et al., 2020), we uniformly sample a certain ratio $r_{tab}$ of the fields, and then replace the corresponding feature with an additional <MASK> feature, which is set as a special feature in the embedding table of ID-based model. The <MASK> feature is not field-specific, but shared by all feature fields to prevent introducing prior knowledge about the masked field (Lin et al., 2023b). Similarly, we denote the index set of masked fields as $\mathcal{I}^{tab}$.
@@ 3.3.2. Data Encoding
(4) $\textbf{w}_{i}$ $=h_{\text{PLM}}(\textbf{x}_{i}^{text}),\;\;\;\hat{\textbf{w}}_{i}=h_{\text{PLM}}(\hat{\textbf{x}}_{i}^{text}),$
@@
Here, $\textbf{w}_{i}=[w_{i,l}]_{l=1}^{L}\in\mathbb{R}^{L\times D_{text}}$ is the set of hidden states from the last layer of the PLM, where $L$ is the number of tokens for $\textbf{x}_{i}^{text}$ and $D_{text}$ is the hidden size of the PLM. Note that $w_{i,1}$ is the [CLS] token vector that represents the overall textual input. $\hat{\textbf{w}}_{i}=[\hat{w}_{i,l}]_{l=1}^{L}\in\mathbb{R}^{L\times D_{text}}$ also satisfies the notations above. $\textbf{v}_{i},\hat{\textbf{v}}_{i}\in\mathbb{R}^{D_{tab}}$ are the representations produced by ID-based model.
3.3.3. Masked Language Modeling
As shown in Figure 2 (Stage 2), the MLM module takes the text-tabular pair $(\hat{\textbf{w}}_{i},\textbf{v}_{i})$ as input, and attempts to recover the masked tokens. We denote the index set of masked tokens as $\mathcal{I}^{text}$. For each masked token with index $l\in\mathcal{I}^{text}$, we concatenate the corresponding token vector $\hat{w}_{i,l}$ with the “reference answer” $\textbf{v}_{i}$, and then feed them through the prediction layer to obtain the estimated distribution:
(5) $q_{i,l}=g_{\text{PLM}}([\hat{w}_{i,l}\oplus\textbf{v}_{i}])\in\mathbb{R}^{V},$
where $\oplus$ is the concatenation operation, $V$ is the vocabulary size, and $g_{\text{PLM}}$ is a two-layer MLP. Finally, similar to common masked language modeling (Devlin et al., 2018), we leverage the cross-entropy loss for pretraining optimization:
(6) $\mathcal{L}^{MLM}_{i}=\frac{1}{|\mathcal{I}^{text}|}\sum\nolimits_{l\in\mathcal{I}^{text}}\operatorname{CrossEntropy}\left(q_{i,l},\ x^{text}_{i,l}\right),$
@@ 3.3.3. Masked Language Modeling
3.3.4. Masked Tabular Modeling
@@
(7) $\textbf{u}_{i}=\text{Softmax}\left(\frac{\widehat{\textbf{v}}_{i}\textbf{Q}\textbf{w}_{i}^{T}}{\sqrt{D_{text}}}\right)\textbf{w}_{i},\;\;\textbf{u}_{i}\in\mathbb{R}^{L}$
where $\textbf{Q}\in\mathbb{R}^{D_{tab}\times D_{text}}$ is the trainable cross-modal attention matrix, and $\sqrt{D_{text}}$ is the scaling factor (Vaswani et al., 2017).
@@
(8) $\textbf{c}_{i,f}$ $=g_{\text{ID}}^{(f)}(\textbf{u}_{i}),\quad\textbf{c}_{i,f}\in\mathbb{R}^{M},$
@@
(9) $\mathcal{L}_{i}^{MTM}=\frac{1}{|\mathcal{I}^{tab}|}\sum\nolimits_{f\in\mathcal{I}^{tab}}\operatorname{CrossEntropy}(p_{i,f},\ x^{tab}_{i,f}),$
@@
However, the loss above is actually impractical and inefficient since it has to calculate the softmax function over the entire feature space in Eq. 8, where $M$ is usually at million level for real-world applications. To this end, we adopt noise contrastive estimation (NCE) (Gutmann and Hyvärinen, 2010; Mikolov et al., 2013; Lin et al., 2023b). NCE transforms a multi-class classification task into a binary classification task, where the model is required to distinguish positive features (i.e., masked features) from noise features. Specifically, for each masked feature $x_{i,f}^{tab}$ of $f$-th field, we sample $K$ noise features from the entire feature space according to their frequency distribution in the training set. Then, we utilize the binary cross-entropy (BCE) loss for MTM optimization:
(10) $\mathcal{L}_{i}^{MTM}=-\frac{1}{|\mathcal{I}^{tab}|}\sum_{f\in\mathcal{I}^{tab}}(\log\sigma(c_{i,f,t})+\sum_{k=1}^{K}\log(1-\sigma(c_{i,f,k}))))$
@@
3.3.5. Instance-level Contrastive Learning (ICL)
In addition to the feature-level alignment through the masked modality modeling, we also introduce contrastive learning to explicitly learn instance-level consistency between two modalities. The contrastive objective draws the representations of matched text-tabular pairs together and pushes apart those non-matched pairs (Radford et al., 2021; Li et al., 2021).
Here, we utilize the [CLS] token vector $w_{i,1}$ to represent the textual input $\textbf{x}_{i}^{text}$. For dimensional consistency, we employ two separate linear layers to project the [CLS] token vector $w_{i,1}$ and tabular representation $\textbf{v}_{i}$ into $d$-dimensional vectors, $z^{text}_{i}$ and $z^{tab}_{i}$, respectively. Next, we adopt InfoNCE (Oord et al., 2018) to compute the ICL loss:
(11) $\begin{split}\mathcal{L}^{ICL}=-\frac{1}{2B}\sum_{i=1}^{B}\Bigg{[}&\log\left(\frac{\text{exp}(\text{sim}(z^{text}_{i},z^{tab}_{i})/\tau)}{\sum_{j}\text{exp}(\text{sim}(z^{text}_{i},z^{tab}_{j})/\tau)}\right)\\ +&\log\left(\frac{\text{exp}(\text{sim}(z^{tab}_{i},z^{text}_{i})/\tau)}{\sum_{j}\text{exp}(\text{sim}(z^{tab}_{i},z^{text}_{j})/\tau)}\right)\Bigg{]}\end{split}$
@@ 3.3.5. Instance-level Contrastive Learning (ICL)
(12) $\mathcal{L}^{pretrain}=\frac{1}{B}\sum\nolimits_{i=1}^{B}\left(\mathcal{L}^{MLM}_{i}+\mathcal{L}^{MTM}_{i}\right)+\mathcal{L}^{ICL}.$
3.4. Adaptive Finetuning
After the pretraining stage, the PLM and ID-based model have learned fine-grained multimodal representations. As depicted in Figure 2 (Stage 3), in this stage, we adaptively finetune the two models jointly on the downstream CTR prediction task with supervised click signal to achieve superior performance.
FLIP places a randomly initialized linear layer on the ID-based model, and another layer upon the PLM, so that these two models can output the estimated probability $\hat{y}_{i}^{\text{ID}}$ and $\hat{y}_{i}^{\text{PLM}}$ respectively.
(13) $\hat{y}^{\text{ID}}_{i}$ $=\sigma(\operatorname{Linear}_{\text{ID}}(\textbf{v}_{i})),$
@@ 3.4. Adaptive Finetuning
(14) $\hat{y}_{i}=\sigma\left(\alpha\times\operatorname{Linear}_{\text{ID}}(\textbf{v}_{i})+(1-\alpha)\times\operatorname{Linear}_{\text{PLM}}(\textbf{w}_{i})\right),$
where $\alpha\in$ is a learnable parameter to adaptively balance the outcomes from two models. To avoid performance collapse mentioned in previous works (Bühlmann, 2012; Schapire, 2013), we apply BCE objectives over the jointly estimated click probability $\hat{y}$, as well as the solely estimated click probability $\hat{y}_{i}^{\text{ID}}$ and $\hat{y}_{i}^{\text{PLM}}$ for model optimization:
(15) $\mathcal{L}^{finetune}=\mathcal{L}_{BCE}(y,\hat{y})+\mathcal{L}_{BCE}(y,\hat{y}^{\text{ID}})+\mathcal{L}_{BCE}(y,\hat{y}^{\text{PLM}})$
@@
4. Experiment
4.1. Experiment Setup
4.1.1. Datasets
We conduct experiments on three real-world public datasets: MovieLens-1M, BookCrossing, GoodReads. All of the selected datasets contain user and item information. The information is unencrypted original text, thus preserving the real semantic information.
  • MovieLens-1M (Harper and Konstan, 2015) is a movie recommendation dataset with user-movie ratings ranging from 1 to 5. Following the previous work (Song et al., 2019; Li et al., 2023a), We consider samples with ratings greater than 3 as positive, samples with ratings less than 3 as negative, and remove samples with ratings equal to 3 (i.e., neutral).
  • BookCrossing (Ziegler et al., 2005) is a book recommendation dataset and possesses user-book ratings ranging from 0 to 10. We consider samples with scores greater than 5 as positive, and the rest as negative.
  • GoodReads (Wan and McAuley, 2018; Wan et al., 2019) is a book recommendation dataset which contains user-book ratings ranging from 1 to 5. We take samples with ratings greater than 3 as positive, and the rest as negative.
Following previous works (Li et al., 2023a, 2022a), we sort all samples in chronological order and take the first 90% samples as the training set and the remaining as the testing set. The training set for the pretraining and finetuning stage are the same (Lin et al., 2023b). All our experimental results are obtained on the testing set. The statistics of the processed datasets are shown in Table 1.
4.1.2. Evaluation Metrics
We use commonly adopted metrics, AUC (Area Under the ROC Curve) and Logloss (binary cross-entropy loss) as the evaluation metrics. Notably, a slightly higher AUC or a lower Logloss (e.g., 0.001) can be considered as a significant improvement in CTR prediction (Lian et al., 2018; Guo et al., 2017; Li et al., 2022b).
4.1.3. Baselines
The baseline methods can be mainly classified into three categories: (1) ID-based models: AFM (Xiao et al., 2017), PNN (Qu et al., 2016), Wide&Deep (Cheng et al., 2016), DCN (Wang et al., 2017), DeepFM (Guo et al., 2017), xDeepFM (Lian et al., 2018), AFN (Cheng et al., 2020), AutoInt (Song et al., 2019) and DCNv2 (Wang et al., 2021), (2) PLM-based models: CTR-BERT (Muhamed et al., 2021), P5 (Geng et al., 2022) and PTab (Liu et al., 2022), (3) ID+PLM models that combine ID-based model and PLM: CTRL (Li et al., 2023a), MoRec (Yuan et al., 2023).
Table 1. Statistics of processed datasets.